# Eaman Jahani

### v4

@@ Abstract
The recent surge in generative AI has led to new models being introduced almost every month. In light of this rapid progression, we pose and address a central question: to what extent must prompts evolve as the capabilities of generative AI models advance? To answer this question, we conducted an online experiment with N = 1,893 participants where each participant was incentivized to write prompts to reproduce a target image as closely as possible in 10 consecutive tries. Each participant was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor, DALL-E 3, or a version of DALL-E 3 with automatic prompt revision. In total, we collected and analyzed over 18,000 prompts and over 300,000 images. We find that task performance was higher for participants using DALL-E 3 than for those using DALL-E 2. This performance gap corresponds to a noticeable difference in the similarity of participants’ images to their target images, and was caused in equal measure by: (1) the increased technical capabilities of DALL-E 3, and (2) endogenous changes in participants’ prompting in response to these increased capabilities. Furthermore, while participants assigned to DALL-E 3 with prompt revision still outperformed those assigned to DALL-E 2, automatic prompt revision reduced the benefits of using DALL-E 3 by 58%. Our results suggest that for generative AI to realize its full impact on the global economy, people, firms, and institutions will need to update their prompts in response to new models.
Recent economic forecasts predict that generative AI could add trillions of dollars to the global economy annually, and there is mounting evidence that this new technology is being integrated into work practices in areas as diverse as business, medicine, and government FactSet (2024); Zhang and Kamel Boulos (2023); Bright et al. (2024). This widespread adoption appears to be warranted, with recent studies finding that giving workers generative AI improves productivity across various professional settings (Brynjolfsson et al., 2023; Dell’Acqua et al., 2023; Noy and Zhang, 2023). A common feature of these studies is that they used static model versions for testing. However, in practice, generative models are continuously updated, with new versions being released as frequently as once per month. People interact with generative AI by providing textual instructions or “prompts” to guide the models’ outputs (Oppenlaender, 2023; Schulhoff et al., 2024; Don-Yehiya et al., 2023; Xie et al., 2023). It remains unclear whether organizations must continually adjust these prompts as the models advance to fully capture their economic and societal advantages. For example, it is now common amongst software companies to have a hundred or more development teams, each with hundreds, if not thousands, of prompts integrated with their code. On the one hand, firms might write these prompts once and subsequently benefit from new models with minimal additional effort. Alternatively, organizations may need to continually adapt their prompts as models improve. In this paper, we investigate whether refining prompts is necessary as models evolve, and how much specialized expertise this refining process requires.
To explore these questions, we conducted a pre-registered online lab experiment. $N=1,893{}$ participants from Prolific were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”). [^1] The main text presents a subset of pre-registered analyses, with additional results and deviations from pre-registration in the SI. The LLM-based prompt revisions in the third condition were not visible to participants during the experiment. [^2] Automatic prompt revision is the default behavior of the DALL-E 3 API. These prompt revisions involved sending each participant’s prompt to GPT-4, which enhanced it by adding descriptive details before being sent to DALL-E 3. We followed the DALL-E 3 API documentation and attempted to disable prompt revision in the second, “DALL-E 3” treatment arm by prepending a system prompt to participants’ prompts. While this drastically reduced the number and extent of LLM-based prompt revisions, it did not completely eliminate them. Therefore, the “DALL-E 3” treatment arm still includes some automatic revisions. We refer to this treatment arm as “DALL-E 3” for brevity and simplicity throughout the paper. Full details on the system prompt’s efficacy are provided in the SI. The task required each participant to make at least 10 attempts at recreating a “target image” as closely as possible by prompting their assigned model. Each participant’s target image was randomly selected from a curated set of 15 images; this set of images was constructed to span realistic use cases of AI-generated images. Throughout the task, participants could view all of their past prompts and generated images alongside the target image. Participants were paid \$4 USD to complete the task, and to incentivize performance, they received \$8 USD if they were in the top 20% of participants, as measured by the similarity of their best attempt to the target image.
Results
To understand how organizations might benefit from model improvements, we first examined whether access to a more advanced generative model enhanced task performance. Comparing the DALL-E 2 and DALL-E 3 treatment arms, we find that providing access to a more advanced generative model improved task performance. Participants using DALL-E 3 produced images that were, on average, $z=0.19$ standard deviations closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$), and the size of this gap did not shrink as participants gained more experience with the task (See Figure 1B; $\beta=0.0010$, $p=0.0227$, see SI for full regression). Figure 1A presents three examples from our data to illustrate the qualitative magnitude of this average treatment effect (ATE), which corresponds to a considerable increase in qualitative similarity to the target image.

### v5

@@ Abstract
As generative AI systems rapidly improve, a key question emerges: How do users keep up—and what happens if they fail to do so. Drawing on theories of dynamic capabilities and IT complements, we examine prompt adaptation—the adjustments users make to their inputs in response to evolving model behavior—as a mechanism that helps determine whether technical advances translate into realized economic value. In a preregistered online experiment with 1,893 participants, who submitted over 18,000 prompts and generated more than 300,000 images, users attempted to replicate a target image in 10 tries using one of three randomly assigned models: DALL-E 2, DALL-E 3, or DALL-E 3 with automated prompt rewriting. We find that users with access to DALL-E 3 achieved higher image similarity than those with DALL-E 2—but only about half of this gain (51%) came from the model itself. The other half (49%) resulted from users adapting their prompts in response to the model’s capabilities. This adaptation emerged across the skill distribution, was driven by trial-and-error, and could not be replicated by automated prompt rewriting, which erased 58% of the performance improvement associated with DALL-E 3. Our findings position prompt adaptation as a dynamic complement to generative AI—and suggest that without it, a substantial share of the economic value created when models advance may go unrealized.
1 Introduction
Generative AI is being integrated into work practices across the economy (Zhang and Kamel Boulos 2023, Bright et al. 2024), yielding notable productivity gains in tasks as diverse as software development, writing, and material discovery (Brynjolfsson et al. 2023, Dell’Acqua et al. 2023, Noy and Zhang 2023, Peng et al. 2023, Yu 2024, Toner-Rodgers 2024). Recent research points to even greater potential ahead, demonstrating advances in automating core scientific processes (Manning et al. 2024), including tasks as complex as chemical research and proving mathematical theorems (Romera-Paredes et al. 2024, Boiko et al. 2023). The adoption of generative AI is also occurring at an unprecedented pace, with recent research showing that approximately 28% of U.S. workers are already using generative AI in their jobs—a rate that significantly outpaces early adoption of personal computers and internet technology at comparable points in their diffusion (Bright et al. 2024, Bick et al. 2024).
As with many other general-purpose technologies, the effectiveness of generative AI depends not only on the technology itself, but on users’ ability to craft inputs that produce high-quality results. To interact with generative AI systems, users provide written instructions—often referred to as prompts—that guide the model’s behavior. These prompts can range from simple commands (e.g., “write a short story about a robot”) to highly detailed specifications tailored to particular outputs (e.g., a series of paragraphs instructing an AI system to implement a complete piece of software). In this way, prompting serves as a complementary skill—one that, like spreadsheet modeling in the early PC era, can determine the productivity impact of a given tool (Brynjolfsson and Hitt 2000).
Prompting has quickly become an area of active research and practice. Scholars have developed taxonomies of prompt engineering techniques (Oppenlaender 2023), documented recurring patterns in prompt construction (Schulhoff et al. 2024), and examined how developers embed prompts into software systems (Liang et al. 2024). Other studies have explored prompting strategies for specific applications, including image generation (Don-Yehiya et al. 2023, Xie et al. 2023) and clinical documentation (Yao et al. 2024). In parallel, practitioners have built prompt libraries, shared tutorials, and developed tools to support prompt design. These developments signal a growing consensus that prompting plays a meaningful role in extracting value from generative AI systems.
Yet despite this consensus, prompting remains understudied as a dynamic practice. Many prompt libraries and tutorials present effective prompts as reusable artifacts. But prompts that work well with one model version may underperform or break entirely with the next (Liang et al. 2024, Meincke et al. 2025). While recent research increasingly views prompting as an adaptive process, empirical evidence remains limited on how these strategies evolve—both as users refine prompts for a single model and as they adjust to model updates—and on how these changes ultimately affect performance. This raises a broader question for individuals and organizations investing in prompting capabilities: Are prompt strategies transferable across model versions, or must they be continually revised to match changing model behavior?
To begin exploring this question, we identify prompt adaptation [^1] We use the phrase “prompt adaptation” to refer specifically to changes in user prompting behavior that arise in response to evolving model capabilities—distinct from the broader practice of prompt engineering, which includes static best practices, libraries, and templates. as a measurable behavioral mechanism through which user-side inputs evolve alongside technical advances. We conceptualize prompt adaptation as a dynamic complement—that is, a user capability that adapts in response to changes in a technological system and is critical to realizing the full economic value of system improvements. In contrast to static complements (e.g., fixed training, prompt templates), dynamic complements emerge through situated use with rapid feedback, respond to model-level change, and may be enabled or suppressed by system design.
To assess the role of prompt adaptation in shaping realized performance—and to separate its contribution from the direct effects of model improvement—we draw on data from a pre-registered online experiment with $1,893{}$ participants. In the study, participants were asked to replicate a target image using prompts submitted to one of three randomly assigned text-to-image models of varying capability: DALL-E 2, DALL-E 3, or DALL-E 3 with an automated large language model (LLM)-based prompt revision. Each participant submitted at least 10 prompts in an effort to reproduce the image as closely as possible, with a large monetary bonus for top performers. By comparing outcomes across arms—and by conducting a posthoc analysis that re-evaluates prompts on alternative models—we estimate the extent to which users adapted their prompts in response to model improvements and how much this adaptation contributed to overall performance.
We find that participants assigned to DALL-E 3 produce significantly more faithful replications than those assigned to DALL-E 2. Importantly, about half of this improvement comes from participants adapting their prompts to exploit the new model’s capabilities—replaying the old DALL-E 2 prompts on DALL-E 3 yields only about half the total improvement. Furthermore, we find that this prompt adaptation is not limited to advanced “prompt engineers”—participants across the full skill distribution benefit from refining their prompts after seeing how the advanced model responds. Finally, we show that an attempt to automate prompt revision via GPT-4 rewriting does not match the performance achieved by manual human adaptation; in fact, it substantially erodes the gains from DALL-E 3. Together, these findings position prompt adaptation as a key mechanism through which users realize the value of rapidly advancing generative AI systems—and as a concrete example of how digital complements must evolve to keep pace with technological change.
In terms of related literature and additional theory, this experiment builds on work in information systems, emphasizing the importance of dynamic, user-driven complements to digital technologies. Research on IT-enabled dynamic capabilities has shown that the value of new systems depends not only on technical infrastructure but on organizations’ ability to reconfigure routines and user behaviors in response to ongoing change (Bharadwaj 2000, Joshi et al. 2010, Teece et al. 1997). Related work on post-adoptive IT use has demonstrated that users often engage only superficially with new systems and that meaningful performance gains tend to emerge only when users experiment with and refine their interaction strategies over time (Jasperson et al. 2005). Recent research on human-AI collaboration further underscores that interface design and task structure shape the degree to which users can learn from and adapt to model behavior (Fügener et al. 2022). And the concept of co-evolution has been introduced to describe how humans and generative AI systems jointly adapt over time, forming interdependent capabilities that neither could realize alone (Böhm and Schedlberger 2023).
We also engage with work on general-purpose technologies, which has long emphasized that the productivity gains from technical advances depend on the development of new human and organizational complements (Brynjolfsson 1993, Brynjolfsson and Hitt 2000, Brynjolfsson et al. 2021, David 1990). We conceptualize prompt adaptation as one such complement—emerging through trial-and-error, accessible across the skill distribution, and potentially hindered when automation misaligns with user intent. In this way, prompt adaptation shapes how and when technical improvements translate into downstream economic value.
Finally, the contribution of the current paper is four-fold. First, we offer direct causal evidence that prompting is not a fixed input but a dynamic capability that co-evolves with model behavior (Böhm and Schedlberger 2023). Second, we show how even non-expert users adapt their prompting in response to model improvements, extending research on IT-enabled dynamic capabilities (Bharadwaj 2000, Joshi et al. 2010, Teece et al. 1997). Third, we document how performance gains emerge through user-initiated trial-and-error rather than static best practices, advancing work on post-adoptive IT use (Jasperson et al. 2005). Finally, we demonstrate that automation intended to reduce user effort can—if misaligned with user intent—undermine the very adaptations that generate value, contributing to ongoing research on human–AI interaction (Fügener et al. 2022, Yao et al. 2024).
The remainder of the paper is organized as follows. We begin by presenting a simple conceptual framework that characterizes how output quality evolves with improvements in model capacity and with users’ corresponding adjustments in effort. We then introduce an experimental design that closely mirrors this framework and describes the data used in our analysis. Next, we present our empirical findings, including a decomposition of the overall effect into components attributable to model improvements versus prompt adaptation, the impact of automated prompt revision, and heterogeneity across user skill levels. Finally, we compare these empirical results to the model’s predictions and conclude by discussing implications for organizations adopting generative AI.
2 Conceptual Framework
As generative AI systems evolve, a user’s ability to adapt prompts for improved models can become an important source of realized performance gains. We develop a stylized conceptual framework to formalize how overall output quality depends on both the model’s capacity and the user’s skill and effort in prompt writing. Our goal is not to create a fully normative model. Rather, we wish to clarify the distinction between improvements directly attributable to the model itself versus those arising from user adaptation as motivation for our experimental design. Additionally, the framework naturally implies predictions about the returns to prompting skill and the heterogeneity of such skill.
Although our empirical setting focuses on image replication, this framework generalizes to other tasks where users interact with generative models—such as text generation, code assistance, or molecule design. In each case, we expect performance to reflect both model capacity and users’ efforts to refine their inputs. We present the core findings here; a complete exposition of the framework with formal proofs is provided in Appendix E.1.
2.1 Notation and Problem Setting
Let $\theta\in(0,1]$ represent the model’s capacity to translate prompts into high-fidelity outputs (for instance, how accurately it captures requested details). Let $s\in(0,1]$ denote a user’s baseline skill in prompt engineering, and let $x\geq 0$ be the effort the user expends on writing and refining prompts. Each unit of $x$ incurs a cost $k\,x$ for some $k>0$, which may reflect the time or cognitive load.